OpenStudy (anonymous):

3x^2+5x+1=0, what is the solution?

OpenStudy (anonymous):

a=3 b=5 c=1 D=b^2-4ac=25-12=13 x1=(-b-sqrtD)/2a=(-5-sqrt13)/6 x2=(-b+sqrtD)/2a=(-5+sqrt13)/6 @Mlitzilplix